# Structure Guided Large Language Model for SQL Generation

### v1

@@ 4.1 Experiment Setup
Model Spider Bird
Exec Acc EM Acc Exec Acc EM Acc
Fine-tuning-based
T5-Base Raffel et al. (2020) 0.5791 0.5722 0.1154 0.1009
T5-Large Raffel et al. (2020) 0.6728 0.6536 0.1970 0.1736
T5-3B Raffel et al. (2020) 0.7440 0.7152 0.2334 0.2068
Structure-learning-based
RAT-SQL + GAP Wang et al. (2019) 0.7505 0.7379 0.2639 0.2431
RASAT + PICARD Qi et al. (2022) 0.8048 0.7537 0.2870 0.2603
[…]
RESDSQL-3B Li et al. (2023a) 0.8173 0.7580 0.3031 0.2569
Graphix-3B + PICARD Li et al. (2023b) 0.8121 0.7563 0.2984 0.2418
Incontext-learning-based
CodeX Chen et al. (2021) 0.7167 0.4905 0.3438 0.3019
PaLM-2 Anil et al. (2023) 0.6985 0.4438 0.2735 0.2543
ChatGPT OpenAI (2023) 0.7394 0.5327 0.3562 0.3041
GPT-4 OpenAI (2023) 0.7665 0.5892 0.4633 0.4255
CoT + ChatGPT Wei et al. (2022) 0.7552 0.5684 0.3665 0.3127
C3-GPT Dong et al. (2023b) 0.8185 0.7036 - -
DIN-SQL + GPT-4 Pourreza and Rafiei (2023) 0.8237 0.7187 0.5079 0.4398
DAIL-SQL + GPT-4 Gao et al. (2023) 0.8449 0.7443 0.5434 0.4581
SGU-SQL + GPT-4 (Ours) 0.8791 0.7679 0.5771 0.4989
@@
[…]
@@
Execution accuracy Model Easy Medium Hard Extra All SGU-SQL + GPT-4 93.5 91.9 845 72.3 87.9 DAIL-SQL + GPT-4 90.6 82.7 75.4 58.2 84.5 DIN-SQL + GPT-4 91.5 78.4 69.4 54.3 82.4 GPT-4 87.3 76.0 63.2 51.9 76.7 CodeX 82.8 67.9 57.7 48.5 73.9 Exact match accuracy Model Easy Medium Hard Extra All SGU-SQL + GPT-4 85.5 83.9 72.69 51.2 76.8 DAIL-SQL + GPT-4 79.5 74.2 65.2 46.2 74.4 DIN-SQL + GPT-4 81.3 67.1 42.7 46.1 71.9 GPT-4 77.9 56.4 37.9 22.5 58.9 CodeX 65.7 42.5 28.7 15.1 49.1
@@ 4.3 Ablation Staudy: Q2
LLM Prompting strategy Exec acc
PaLM-2 + Few-shot Prompting 0.6985
+ CoT Prompting 0.6873
+ SGU-SQL 0.7395
CodeX + Few-shot Prompting 0.7167
+ CoT Prompting 0.7198
+ SGU-SQL 0.7418
ChatGPT + Few-shot Prompting 0.73947
+ CoT Prompting 0.7552
+ SGU-SQL 0.7846
GPT-4 + Few-shot Prompting 0.7665
+ CoT Prompting 0.7834
+ SGU-SQL 0.8791

### v2

@@ 4.1 Experiment Setup
{adjustwidth}
-0.5 cm-0.5 cm Model Spider Bird Exec Acc EM Acc Exec Acc EM Acc Fine-tuning-based T5-Base Raffel et al. (2020) 0.5791 0.5722 0.1154 0.1009 T5-Large Raffel et al. (2020) 0.6728 0.6536 0.1970 0.1736 T5-3B Raffel et al. (2020) 0.7440 0.7152 0.2334 0.2068 Structure-learning-based RAT-SQL + GAP Wang et al. (2019) 0.7505 0.7379 0.2639 0.2431 RASAT + PICARD Qi et al. (2022) 0.8048 0.7537 0.2870 0.2603 S2SQL + ELECTRA Hui et al. (2022) 0.7643 0.7285 - - RESDSQL-3B Li et al. (2023a) 0.8173 0.7580 0.3031 0.2569 Graphix-3B + PICARD Li et al. (2023b) 0.8121 0.7563 0.2984 0.2418 Incontext-learning-based CodeX Chen et al. (2021) 0.7167 0.4905 0.3438 0.3019 PaLM-2 Anil et al. (2023) 0.6985 0.4438 0.2735 0.2543 ChatGPT OpenAI (2023) 0.7394 0.5327 0.3562 0.3041 GPT-4 OpenAI (2023) 0.7665 0.5892 0.4633 0.4255 CoT + ChatGPT Wei et al. (2022) 0.7552 0.5684 0.3665 0.3127 C3-GPT Dong et al. (2023b) 0.8185 0.7036 - - DIN-SQL + GPT-4 Pourreza and Rafiei (2023) 0.8237 0.7187 0.5079 0.4398 DAIL-SQL + GPT-4 Gao et al. (2023) 0.8449 0.7443 0.5434 0.4581 SGU-SQL + GPT-4 (Ours) 0.8791 0.7679 0.5771 0.4989
@@
Baselines To valid the effectiveness of SGU-SQL, we compare it with the state-of-art baselines according to our best knowledge. Following the taxonomy aforementioned in Section 2, we divide all baselines into three categories: $(i)$ Fine-tuning-based: T5-base Raffel et al. (2020), T5-large Raffel et al. (2020), T5-3B Raffel et al. (2020); $(ii)$ structure-learning-based: RAT-SQL Wang et al. (2019), RASAT Qi et al. (2022), S2SQL Hui et al. (2022) ,RESDSQL Li et al. (2023a),GRAPHIX Li et al. (2023b); and $(iii)$ incontext-learning-based: PaLM-2 Anil et al. (2023), CodeX Chen et al. (2021), ChatGPT OpenAI (2023), GPT-4 OpenAI (2023), C3-GPT Dong et al. (2023b), DIN-SQL Pourreza and Rafiei (2023), DAIL-SQL Gao et al. (2023).
@@
{tblr} cells = c, row1 = Silver, row8 = Silver, cell11 = c=6, cell81 = c=6, hlines, Execution accuracy &
Model Easy Medium Hard Extra All
SGU-SQL + GPT-4 93.5 91.9 845 72.3 87.9
DAIL-SQL + GPT-4 90.6 82.7 75.4 58.2 84.5
DIN-SQL + GPT-4 91.5 78.4 69.4 54.382.4
GPT-4 87.3 76.0 63.2 51.9 76.7
CodeX 82.8 67.9 57.7 48.5 73.9
Exact match accuracy
Model Easy Medium Hard Extra All
SGU-SQL + GPT-4 85.5 83.9 72.69 51.2 76.8
DAIL-SQL + GPT-4 79.574.265.246.274.4
DIN-SQL + GPT-4 81.3 67.1 42.7 46.1 71.9
GPT-4 77.9 56.4 37.9 22.5 58.9
CodeX 65.7 42.5 28.7 15.1 49.1
@@ 4.3 Ablation Staudy: Q2
{adjustwidth}
-0. cm-0. cm LLM Prompting strategy Exec acc PaLM-2 + Few-shot Prompting 0.6985 + CoT Prompting 0.6873 + SGU-SQL 0.7395 CodeX + Few-shot Prompting 0.7167 + CoT Prompting 0.7198 + SGU-SQL 0.7418 ChatGPT + Few-shot Prompting 0.73947 + CoT Prompting 0.7552 + SGU-SQL 0.7846 GPT-4 + Few-shot Prompting 0.7665 + CoT Prompting 0.7834 + SGU-SQL 0.8791